### v1

# Multimodal Fact-Checking: An Agent-based Approach

Abstract
The rapid spread of multimodal misinformation poses a growing challenge for automated fact-checking systems. Existing approaches, including large vision–language models (LVLMs) and deep multimodal fusion methods, often fall short due to limited reasoning and shallow evidence utilization. A key bottleneck is the lack of dedicated datasets that provide complete real-world multimodal misinformation instances accompanied by annotated reasoning processes and verifiable evidence. To address this limitation, we introduce RW-Post, a high-quality, explainable dataset for real-world multimodal fact-checking. RW-Post aligns real-world multimodal claims with their original social media posts, preserving the rich contextual information in which the claims are made. In addition, the dataset includes detailed reasoning and explicitly linked evidence, which are derived from human-written fact-checking articles via large language model–assisted extraction pipeline, enabling comprehensive verification and explanation. Building upon RW-Post, we propose AgentFact, an agent-based multimodal fact-checking framework designed to emulate the human verification workflow. AgentFact consists of five specialized agents that collaboratively handle key fact-checking subtasks, including strategy planning, high-quality evidence retrieval, visual analysis, reasoning, and explanation generation. These agents are orchestrated through an iterative workflow that alternates between evidence searching, task-aware evidence filtering and reasoning, facilitating strategic decision-making and systematic evidence analysis. Extensive experimental results demonstrate that the synergy between RW-Post and AgentFact substantially improves both the accuracy and interpretability of multimodal fact-checking.
I Introduction
The proliferation of social media has intensified the spread of persuasive textual and visual misinformation, posing serious risks to public health and behavior [borges2022infodemics, wsj2025falseTariffHeadline, ARCURI2023106130]. Advances in generative AI further exacerbate this threat by enabling realistic, multimodal fake content [berman2024aimisinformation, reuters2025duterteDisinformation]. Traditional binary fact-checking is inadequate, as it lacks rationales and evidence that foster public trust [blanchar2024trump, Porter2021]. Narrow tasks like out-of-context (OOC) detection [qi2024sniffer] fall short in addressing real-world complexities. To enhance transparency and reliability, explainable multimodal fact-checking has become a crucial research focus.
Recent work has explored LLM-based methods using techniques such as chain-of-thought prompting, claim decomposition, question-guided prompting, and knowledge distillation [PanQACheck23, leite2023detecting, zhang2023towards, kareem2023fighting]. LVLM-based frameworks have extended these techniques to multimodal settings by incorporating image search and external retrieval, evolving into agent-collaborative systems [wang2024mmidr, xuan2024lemma, tonglet2024image, qi2024sniffer, DuICML24, braun2025defame].
While these efforts move toward more interpretable fact-checking, they still oversimplify the reasoning required for fact-checking and rely heavily on LLMs’ direct generation based on superficially retrieved evidence (e.g., retrieved with claim or image queries). Besides, a major bottleneck remains: no existing dataset offers both realistic multimodal claims and fully aligned, explanation-level annotations with source-grounded evidence. This limits the evaluation and development of reasoning-aware fact-checkers.
To address this gap, we propose an integrated solution that combines a new dataset with a collaborative fact-checking framework composed of specialized agents. We introduce the Real-World Fact-Checking Dataset (RW-Post), which pairs Snopes claims with multimodal (text-image) posts, enriched with fine-grained reasoning, evidence descriptions, and verified source links, allowing agent-based systems to emulate human-like fact-checking with enhanced reasoning and evidence grounding. Complementing the dataset, we present AgentFact, a collaborative agent-based framework that emulates the human fact-checking workflow. AgentFact decomposes verification into a series of specialized modules—Strategy Planning, Visual Retrieval and Analysis, Text Evidence Retrieval, Plan Guided Reasoning, and Explanation Generation—each designed to handle a key subtask. Key strengths of AgentFact include: (1) systematic reasoning grounded in real-world fact-checking practices, (2) modular agents with task-specific expertise, and (3) task-aware utilization of both textual and visual evidence. Our main contributions are:
1.
We release RW-Post, a high-quality, real-world dataset for explainable multimodal fact-checking, offering comprehensive annotations of claims, posts, reasoning steps, and source-grounded evidence.
2.
We propose a dataset collection and annotation pipeline based on existing fact-checking website articles, enabling the construction of a reliable and explainable multimodal dataset for emerging misinformation.
3.
We propose AgentFact, an agent-based framework that models the human verification process via modular, strategy-driven reasoning, supporting robust fact-checking across diverse misinformation types. Experiments on 4 benchmark datasets demonstrate its superiority, especially in multimodal scenario and explanation tasks. The dataset and code will be released at https://github.com/xudanni0927/AgentFact.
II Related Work
II-A Misinformation and OOC Detection Datasets
Misinformation datasets can be broadly grouped into three categories: (1) collected datasets, (2) synthetic datasets, and (3) OOC-specific datasets, as summarized in Table I.
Collected datasets.
Early datasets compile verified claims from fact-checking sites such as Weibo, PolitiFact, and Snopes (e.g., Weibo[nan2021mdfend], FakeNewsNet [shu2020fakenewsnet], LIAR-RAW [yang-etal-2022-coarse], RAWFC [yang-etal-2022-coarse], MuMiN [NielsenMcConville2022] , MR2 [hu2023mr2]). To balance labels, real claims are sometimes sourced from trusted outlets (e.g., Xinhua News). Most are text-only and suffer from noisy annotations or insufficient context. For example, Twitter, Factify 2, and MuMiN datasets match tweets to claims via keyword or cosine similarity, which fails to preserve actual post context. Larger-scale corpora like Fakeddit [nakamura2020r] rely on weak labeling from platform tags, resulting in noisy labels. AVERITEC [schlichtkrull2023averitec] highlights the importance of claim context and employs a multi-step human annotation procedure, producing context-independent and high-quality explanations by having annotators inspecting fact-checking articles. However, AVERITEC remains text-only and does not address multimodal claims.
Several recent efforts attempt to enrich multimodal fact-checking resources. CLAIMREVIEW2024+ (CLAIMREVIEW+) recently curated 140 manually linked images. A few datasets also provide evidence or explanations. For example, Mocheg heuristically extracts rationales from Snopes articles, but unstable heuristics and missing original posts limit the reliability of its annotations.
Synthetic and OOC datasets.
Synthetic datasets enhance diversity and scale through human or model-generated claims. FEVER, FEVEROUS, and HOVER mutate/paraphrase Wikipedia content to create verified claims with increasing reasoning depth. Recent work leverages LLMs to generate claims (e.g., Multi-News-Fact-Checking (Multi-News FC) [chen2024metasumperceiver] using LLaMA2), but these often rely only on textual entailment, neglecting multimodal cues. OOC-specific datasets focus on mismatches between real images and misleading text, using (1) random mismatches (e.g., MAIM, COSMOS [jaiswal2017multimedia, aneja2023cosmos]), (2) semantically controlled mismatches (e.g., NewsCLIPpings [luo2021newsclippings]), or (3) entity substitution [papadopoulos2023synthetic, muller2020multimodal]. MMFakeBench introduces both multimodal and multi-type falsified posts (OOC, visual and textual edits) via LLM/VLM generation and recombination. Despite progress, most synthetic or OOC datasets lack realism, interpretability, and explicit explanations.
These limitations highlight the need for a more reliable and realistic multimodal fact-checking dataset with accurate context and high-quality explanations. To address this gap, we introduce RW-Post—a real-world dataset of text-image misinformation, featuring authentic image and post context, structured rationales, and rationale-aligned evidence.
II-B Misinformation and OOC detection methods
Approaches to misinformation and out-of-context (OOC) detection can be classified along two axes: (1) internal estimation vs. external retrieval, and (2) feature-based vs. LLM/LVLM-based reasoning. Feature-based internal methods rely on writing style, propagation patterns, or image-text consistency [przybyla2020capturing, shu2020hierarchical, zhou2020multimodal, luo2021newsclippings], but lack external grounding and robustness. Feature-based external approaches retrieve related content (e.g., captions, images) and compare via semantic or entity-level similarity [aneja2023cosmos, muller2020multimodal, Papadopoulos_2025_WACV, abdelnabi2022open], though often lacking deeper reasoning. LLM/LVLM-based internal reasoning uses prompting and knowledge distillation [PanQACheck23, leite2023detecting, zhang2023towards], but is limited by internal knowledge and modality coverage. LLM/LVLM-based external methods leverage instruction-tuned models or agent-based designs with retrieval [qi2024sniffer, DuICML24, xuan2024lemma, braun2025defame], enhancing interpretability but still falling short on multi-step evidence synthesis.
However, existing methods often overestimate the reasoning capabilities of large models while underestimating the complexity of fact-checking. Authenticity assessment goes beyond retrieving related captions or images—it demands high-level planning, investigative effort, and both local and global evidence analysis and reasoning, as practiced by human fact-checkers. This is a systematic and cognitively intensive process, where each subtask requires substantial effort. To address these challenges, we propose a collaborative multi-agent framework that decomposes multimodal fact-checking into specialized subtasks, each handled by dedicated agents, designed to emulate real-world verification workflows.
| Dataset | Modality | Time | Size | ||||
| Collected Fact-checking Datasets | |||||||
| T, I | ✗ | ✓ | ✗ | ✗ | Var. | 13k+ | |
| T, I | ✗ | ✓ | ✗ | ✗ | 12–16 | 9k+ | |
| FakeNewsNet | T, I | ✓ | ✓ | ✗ | ✗ | Var. | 19k+ |
| Fakeddit | T, I | ✗ | ✓ | ✗ | ✗ | 08–19 | 1M+ |
| Weibo21 | T, I | ✗ | ✓ | ✗ | ✗ | 14–21 | 9k+ |
| MuMiN | T, I | ✓ | ✓ | ✗ | ✗ | N/A | 21M+ |
| RAWFC | T | ✓ | ✗ | ✓ | ✓ | 07–16 | 2k |
| LIAR-RAW | T | ✓ | ✗ | ✓ | ✓ | 07–16 | 12k |
| AVERITEC | T | ✓ | ✓ | ✓ | ✓ | –21 | 4.57k |
| Mocheg | T, I | ✓ | ✗ | ✓ | ✓ | N/A | 15k |
| Factify 2 | T, I | ✓ | ✓ | ✗ | ✓ | N/A | 50k |
| MR2 | T, I | ✓ | ✓ | ✗ | ✓ | –22 | 14k+ |
| CLAIMREVIEW+ | T, I | ✓ | ✗ | ✗ | ✗ | 23–25 | 0.3k |
| RW-Post (ours) | T, I | ✓ | ✓ | ✓ | ✓ | 17–24 | 1.77k |
| Synthetic Fact-checking Datasets | |||||||
| HOVER | T | ✓ | ✗ | ✗ | ✓ | N/A | 26k |
| FEVEROUS | T | ✓ | ✗ | ✗ | ✓ | N/A | 87k |
| Multi-News FC | T, I | ✓ | ✗ | ✗ | ✓ | –19 | 1291k |
| OOC-specific Datasets | |||||||
| MMFakeBench | T, I | ✗ | ✓ | ✗ | ✗ | N/A | 11k |
| MAIM | T, I | ✗ | ✓ | ✗ | ✗ | N/A | 239k |
| COSMOS | T, I | ✗ | ✓ | ✗ | ✗ | N/A | 453k |
| NewsCLIPpings | T, I | ✗ | ✓ | ✗ | ✗ | N/A | 988k |
III Multi-modal Misinformation Dataset with MM Claim, Post and Rationale
We introduce RW-Post, a high-quality and explainable multimodal fact-checking dataset designed to reflect the complexity of real-world verification scenarios. Each claim in RW-Post is linked to its original context post, reasoning and itemized key points, and precisely aligned evidence. Unlike previous datasets that often include loosely relevant documents, RW-Post emphasizes targeted evidence that directly supports factual reasoning, aligning with the rigor of professional fact-checking. Following the design of benchmarks widely used in large model and fact-checking evaluations [yu2024mm, hendryckstest2021, liu2024mmfakebench], our dataset is structured with both a development set and a test set, with a larger portion allocated to testing. The data are split at a 1:4.29 ratio, yielding 335 development samples and 1,436 test samples. The development set is used for framework development and refinement, while the test set enables fair model comparison. To build the dataset, we collected 13,000 Snopes articles (2017–2024), followed by strict filtering and quality control. For structured extraction of claims, referenced images, posts, reasoning, and evidence from HTML, we apply a reformatting strategy that converts webpages into a text–with-tagged-URL sequence. This enables LLM to better interpret the contextual meaning of each URL and accurately extract relevant content. The pipeline is illustrated in Figure 1.
III-A Data collection and filtering
Data collection. We collect fact-checking articles from Snopes, which follows professional verification standards and covers misinformation from various sources and over ten domains, ensuring both reliability and diversity. Specifically, we extract the following key fields: article URL, category, headline, claim, truthfulness label and full HTML article content.
Text and image size filtering. We filter out articles that are either too short or too long. Extremely short articles—for example, those containing only a single sentence and a link—typically lack sufficient information for verification. Conversely, overly long articles often include excessive background details that may hinder authenticity assessment. Ultimately, we retain articles containing between 90 and 2,000 words. Images smaller than 200200 pixels are also excluded.
III-B Dataset annotation
1. Relabeling Truthfulness Label. The Snopes website provides a wide range of fine-grained truthfulness labels (over ten), but the boundaries between some of them are ambiguous. For example, categories like MISCAPTION and FALSE often overlap. Additionally, the label distribution is highly imbalanced. To enable more reliable classification and analysis, we consolidate the filtered articles into five high-level categories based on authenticity: True, False, Unproven, Mixture, and Outdated—by relabeling the original Snopes labels accordingly. The mapping from Snopes’s original labels to our consolidated categories is shown below. Note that we exclude label types not directly related to factual accuracy, such as satire.
| Label | Original Labels |
| False | FALSE, MISCAPTION, MISCAPTIONED, FAKE, LEGEND, SCAM, MISATTRIBUTED |
| Mixture | MOSTLY FALSE, MIXTURE, MOSTLY TRUE |
| Outdated | OUTDATED |
| True | TRUE, CORRECT ATTRIBUTION, LEGIT |
| Unproven | UNFOUNDED, UNPROVEN, RESEARCH IN PROGRESS |
2. Reformating article. Inspired by Flamingo [flamingo-2022], which demonstrates the effectiveness of interleaving text and media for contextual grounding, we reformat each article into a text-with-tagged-URLs structure. This format is critical for accurately extracting claim-aligned posts and rationale-aligned evidence. As shown in Figure 1, retaining both the tag and its original position preserves the contextual meaning of each URL. Specifically, we categorize all URLs in the HTML as image, video, or webpage links based on file extensions, then tag and embed them within the text accordingly. The result is an article where text is seamlessly interleaved with tagged multimodal content.
3. Extracting Explanation. While fact-checking articles typically include essential reasoning and evidence, this information is often unstructured and intermingled. To address this, we design a ground truth generation pipeline that leverages the text-with-URL-tags format and LLM to extract claim-aligned posts, images, fact-checking rationales, and rationale-aligned evidence. (1) We use GPT-4o to automatically extract post content from articles. The interleaved text-URL format enables GPT-4o to effectively identify high-quality multimodal posts, incorporating images and videos when available. (2) We prompt GPT to summarize each text-URL article’s fact-checking rationale into a concise reasoning logic overview and a list of key points. (3) For each key point, we further prompt GPT to extract supporting evidence, providing a short evidence description and the corresponding evidence links (e.g., web, image, or video URLs).
III-C Quality Control
To ensure the quality of the extracted posts, we filtered out non-multimodal claims—i.e., claims that do not require image support for verification—as well as images containing watermarks indicating fact-checking results, to prevent label leakage into detection models. This filtering process combines GPT-based prompting with human verification. Specifically:
1.
To determine whether a claim requires image context for verification, we designed a prompt specifying that an image is necessary if it serves as direct evidence, enhances the claim’s credibility, or significantly influences the perceived truthfulness of the information. To improve quality, GPT was asked to select from three options—yes, no, and uncertain—rather than making a binary decision. A sample of 70 outputs labeled as yes or no was manually reviewed by human annotator, and all were found to be consistent with GPT’s predictions. Claims labeled as no or uncertain were excluded from the dataset to ensure reliability.
2.
To identify whether an image contains fact-checking-related watermarks, GPT was prompted to analyze the visual content for keywords such as fake, false, or misleading. In a manual evaluation of 50 samples, GPT achieved an accuracy of 98%.
The full prompt templates for both tasks are provided in the Appendix -A.
III-D Human validation
To ensure that the posts, images, rationales, and evidence extracted by LLMs are reliable and each sample is really text-image claim, we conduct a human validation study.
Validation of Post and Image Extraction
Annotators judge whether the post is correctly extracted, and whether the image is necessary or relevant for understanding the claim, by manually comparing the original fact-checking article with the extracted post and image. Each sample was independently annotated by two reviewers, and all disagreements were resolved through discussion until consensus was reached.
For post-link extraction, annotations distinguish between correct extractions, missing extractions, incorrect extractions, and false positives, enabling a fine-grained analysis of the LLM-annotated result. On a manually reviewed set of 100 Snopes articles, the system correctly extracted post links in 76.39% of link-containing cases, with errors primarily arising from missed extractions rather than incorrect matches, and produced false positives in 14.29% of link-absent articles.
For image annotations, claims are categorized based on (i) whether the image is semantically related to the claim and (ii) whether the image content or authenticity affects the claim’s veracity, resulting in three categories: necessary, relevant, and irrelevant. The results show that images are evidentially necessary in 87% of claims, while 5% involve related but non-decisive images, and only 8% correspond to erroneous cases with irrelevant image usage. Overall, these distributions indicate high annotation quality and support the reliability of the dataset for evaluating multimodal fact-checking. Figure 2 illustrates two examples validated by human annotators.
In summary, the correctness of the text–image claims and the role of images are strongly validated, while the post-extraction capability of the LLM is reasonably accurate but still limited in completeness and can be further improved.
Validation of Reasoning and Evidence
Human evaluation of the extracted reasoning, key points, and supporting evidence follows the same procedure used for evaluating the fact-checking models’ outputs, and is therefore described in our main experimental section V-D.
III-E Statistics
Summary statistics are shown in Figure 3. The RW-Post dataset spans 10 domains, with Fauxtography (831), Politics (384), Entertainment (107), Junk News (99), and Viral Phenomena (64) being the most frequent categories.
The main dataset consists of 1.77k main instances in three veracity types: False, True, and Unproven (Figure 3a). We also include two less common but important categories, Mixture and Outdated, which capture more complex or time-sensitive claims. All Mixture and Outdated cases are grouped into a separate set of 218 challenging samples.
Most articles contain diverse modalities of evidence such as text, image, and multimodal combinations (Figure 3b), and mostly contain 3–8 pieces of evidence (Figure 3c). Categories like History, Health, and Viral Phenomena show higher median evidence counts (Figure 3d), indicating that human fact-checkers might conduct more in-depth investigations for these topics.
IV Method
Overview
We decompose the fact-checking problem into five independent sub-tasks and design a dedicated agent for each of them. Building on these components, we develop a fact-checking pipeline, termed the AgentFact framework, which integrates the five agents into a multi-round evidence retrieval, filtering and reasoning process to achieve high-quality fact checking. Unlike prior approaches that treat fact checking as a single autonomous task—often reduced to question answering with the help of an external search engine—AgentFact models fact checking as a collaborative workflow in which each agent is specialized and explicitly designed with fact-checking knowledge. Our approach not only determines the veracity of claims but also produces detailed, evidence-grounded justifications for its decisions. We next introduce our explainable fact-checking problem formulation, agent designs, and overall workflow.
IV-A Problem Definition
We present a structured assessment of the authenticity and explanation in multimodal fact-checking. The input is a multimodal claim, represented as a tuple , where:
is the textual Claim,
is the associated Post Context, composed of post text and image .
The objective is to assess the veracity of the multimodal claim and provide supporting reasoning and evidence. The output is defined as a tuple:
where:
represents the veracity assessment result (a discrete label or a continuous falsity score),
is the complete set of evidence items, including textual and/or visual sources,
is a set of key reasoning points , each aligned with a subset of supporting evidence ,
is a textual reasoning logic that summarizes the key reasoning points,
is a confidence score that represents the certainty of the prediction.
This formulation enables fine-grained multimodal claim assessment through rigorous reasoning–evidence alignment.
IV-B Agents for Fact-Checking
As illustrated in Figure 4, the framework consists of five specialized agents: 1) Strategy Planning Agent (Agent-SP) , 2) Text Evidence Retrieval and Validation Agent (Agent-TR), 3) Image Retrieval and Analysis Agent (Agent-IR), 4) Reasoning Agent (Agent-R), 5) Explanation Generation Agent (Agent-EG). Each agent is designed with task-specific knowledge to mimic distinct components of real-world fact-checking workflows, and is deployed using prompt-engineering techniques with tool interaction. The simplified prompts are provided in the Appendix -B.
IV-B1 Strategy Planning Agent (Agent-SP)
In real-world fact-checking organizations, different types of misinformation require different verification strategies. To reflect this, we employ an LLM to generate tailored, strategy-guided fact-checking plans for each claim. Specifically, Agent-SP draws inspiration from example reasoning strategies such as Divide and Conquer, Chain of Evidence, Cross-Verification, Primary Source Verification, and Temporal Consistency Check. It takes as input the claim (), the associated post (), and—when applicable—the reasoning trace () from a previous, unsuccessful attempt. The output comprises three components: (1) a structured sequence of reasoning steps (); (2) a validation list (), which contains original statements that require verification, often used as direct search queries; (3) a search list (), which identifies auxiliary information needed for carrying out the current fact-checking plan.
IV-B2 Text Evidence Retrieval and Validation Agent (Agent-TR)
Evidence retrieval is crucial for fact-checking, as irrelevant or misleading evidence can easily lead to undecided or incorrect conclusions. To provide comprehensive and context-rich evidence, our retrieval module operates through four coordinated sub-steps:
1.
Query Generation: Constructs a set of search-engine-optimized queries () based on the search list .
2.
Search: Uses the Serper API111https://serper.dev/ to retrieve evidence based on each query in as well as the queries in the validation list .
3.
Query-Guided Filtering and Summarization: Filters and condenses the retrieved text. Only information directly relevant to the input query is retained and summarized, ensuring compact yet informative evidence.
4.
Source Reliability Evaluation: Our preliminary experiments indicate that, during fact-checking reasoning, LLMs often overlook source reliability, which can lead to incorrect conclusions when models are exposed to deceptive or satirical information sources. To address this issue, we incorporate an additional source reliability evaluation module that classifies each source into four categories: reliable, unreliable, satire, and unsure, producing reliability-annotated evidence denoted as .
IV-B3 Image Retrieval and Analysis Agent (Agent-IR)
Among image-related misinformation types, image manipulation and miscaptioning are the two most common strategies. To detect such cues, we design an Image Retrieval and Analysis Agent that identifies image-based inconsistencies. Notably, retrieved images play different evidential roles in fact-checking depending on their relationship to the input image and the claimed event. Near-duplicate images provide strong evidence, images of the same event provide semantic consistency but require time alignment, and unrelated images tend to introduce noise. To account for these nuances, we define a multi-step pipeline for Agent-IR:
1.
Reverse Image Search: The Google Cloud Vision API is used to retrieve a set of visually similar images to the input image, denoted as . For each retrieved image, we also collect the associated textual content from its source webpage, forming a corresponding text set .
2.
Image Matching Analysis: Retrieved images are categorized by an LLM into three groups: (1) Potentially From the Same Source. Images that are nearly identical to the input image in terms of composition, content, and configuration. (2) Same Event, Different Content. Images that depict the same real-world event as the input image but differ in viewpoint, timing, or framing. (3) No Close Relationship. Images that are visually or semantically unrelated to the input image or the claimed event. The retrieved images in categories (1) and (2) are denoted as , and their associated textual content are denoted as .
3.
Visual Tampering Detection: The post image is compared with to estimate the likelihood of visual manipulation. The output is a tampering score .
4.
Miscaption Detection: The input claim and post text are compared with to assess the likelihood of miscaptioning .
5.
Deepfake Detection: Deepfake detection is performed using the DIRE model [wang2023dire], which returns a deepfake probability score .
IV-B4 Reasoning Agent (Agent-R)
The reasoning agent is provided with the claim (), the post text (), and the validation plan (). It follows the prescribed reasoning steps, evaluating the relevant evidence at each step. Before the reasoning process begins, Agent-R is prompted to rephrase the input claim to improve focus and prevent the model from drifting away from the core verification target. The output () of Agent-R consists of three components: Reasoning Results — step-wise analysis, supporting evidence descriptions, and utility assessments; Veracity Label — the predicted truthfulness of the claim (e.g., true, false, unproven); Confidence Level — a score from 1 to 5 reflecting the model’s certainty.
IV-B5 Explanation Generation Agent (Agent-EG)
The Explanation Generation Agent takes as input the reasoning outputs (), together with the supporting textual and visual evidence, and produces a structured explanation. The output format follows the schema defined in the problem formulation.
We next present how these agents operate together in the AgentFact workflow.
IV-C Workflow
AgentFact employs a workflow that integrates the five agents described above to achieve effective evidence selection and autonomous fact checking. It follows a self-updating planning and a multi-round evidence process paradigm composed of the following stages, illustrated in Figure 5:
Step 1: Plan Generation. The Agent-SP generates an initial fact-checking plan based on the input claim and context. It outlines the reasoning logic and produces two key components: a validation list that specifies the items to be verified, and a search list that enumerates related topics to investigate.
Step 2: Text Evidence Retrieval. The Agent-TR sequentially processes both lists. For each item, it initiates a retrieval round that gathers relevant textual content, applies coarse filtering, and performs query-guided summarization to condense the information. This multi-round process results in an evolving reliable evidence pool .
Step 2*: Visual Evidence Retrieval and Analysis. The Agent-IR operates in parallel, focusing on image-based misinformation cues such as tampering, miscaptioning, and deepfake manipulation. Its outputs are integrated into the accumulated reasoning stage to support a holistic verification process.
Step 3–4: Dual Reasoning. Large amount of evidence and noisy evidence often interrupts the fact-checking reasoning process. To address this, we propose a dual-reasoning mechanism that deploys two reasoning agents to achieve two complementary goals: (1) filtering evidence to identify reasoning-relevant information, and (2) performing enhanced reasoning.
Step 3: Temporary Reasoning Agent (Agent-R-T). This agent is triggered when the reliable evidence pool contains at least a threshold number of items (default = 3). A subset of evidence () are extracted from and provided to Agent-R-T. In this stage, the goal is not to produce a final veracity judgment; instead, Agent-R-T also identifies and returns the evidence set () that supports its reasoning (), thereby reducing noisy evidence and enhancing subsequent reasoning. The selected evidence forms the useful evidence pool .
Step 4: Accumulated Reasoning Agent (Agent-R-A): This agent is triggered whenever new evidence enters . Agent-R-A integrates all evidence in and performs accumulated reasoning to produce a more evidence-grounded judgment .
Step 5: Explanation or Plan Regeneration: During each reasoning cycle, the reasoning agents assess whether sufficient evidence has been gathered to reach a confident conclusion. If the predicted label is not unproven and the confidence score exceeds a threshold (default = 4), or if the number of reasoning steps reaches a limit (default = 12), all reasoning outputs and associated evidence are passed to the Agent-EG. If these conditions are not met, another round of retrieval and reasoning is initiated. If all items in the validation and search lists have been exhausted without reaching a confident conclusion, the strategy planning agent is reactivated to generate a revised plan based on the current intermediate reasoning state.
V Experiments
We evaluate fact-checking performance across two primary tasks: claim classification and explanation generation, which measures the quality of the reasoning (including key points) and the evidence.
V-A Datasets and Comparison Methods
Datasets. Besides the proposed RW-Post dataset, we evaluate AgentFact on recent text-image misinformation detection datasets: Mocheg[mocheg23] and CLAIMREVIEW2024+ [braun2025defame], and out-of-context (OOC)-specific dataset: NewsCLIPpings [luo2021newsclippings]. Due to cost constraints, we use 1,000 sampled instances from each dataset except RW-Post and CLAIMREVIEW2024+, for which we evaluate on the full available test set.
Models. We compare AgentFact against both traditional feature-based multimodal fact-checking methods: CCN [abdelnabi2022open], and large-model-based approaches including LEMMA [xuan2024lemma], DEFAME [braun2025defame], and Sniffer [qi2024sniffer]. Without performing any additional fine-tuning, we conducted pure inference (zero-shot evaluation) on the same selected datasets, aligning our evaluation protocol with other comparative models for a fair assessment.
V-B Implementation Details
Model and API Settings: For fair comparison across all methods, we standardize the use of underlying components where applicable. All methods—except for the CCN model—utilize the GPT-4o-mini API as the large language model (LLM). For approaches that require external web search (including LEMMA, DEFAME, and AgentFact), we employ the Serper API as a unified and consistent search backend.
For Sniffer model, we adopt the publicly released pretrained checkpoints222Sniffer: https://github.com/MischaQI/Sniffer?tab=readme-ov-file provided by the original authors, trained on the NewsCLIPpings dataset. We conduct inference-only evaluation without any additional fine-tuning, following the official implementation for feature extraction and reasoning.
For CCN model, we report the results as provided in the original papers, which were obtained by training and evaluating the models on the full NewsCLIPpings dataset. When directly applying the released checkpoints333CCN: https://github.com/S-Abdelnabi/OoC-multi-modal-fc to the NewsCLIPpings subset used in our experiments, we observed substantially degraded performance, likely due to differences in data splits and preprocessing. As retraining on the exact dataset split requires access to the full dataset, which is currently unavailable to us, we defer a fair re-evaluation to future work.
Computing Resources: All experiments on large model-based methods were conducted on a single PC without any GPU support.
Data Leakage Prevention: To ensure fair evaluation and prevent data leakage, we apply strict filtering to all retrieved evidence used before final reasoning for all external evidence based methods. Specifically, we exclude: (1) any evidence retrieved from Snopes or other major fact-checking websites (see details in Appendix -C), and (2) any evidence published after the original fact-checking date of the claim.
V-C Performance of Veracity
V-C1 Metrics
We evaluate veracity prediction using two classification settings: binary and three-class classification.
For the out-of-context (OOC) task, we conduct binary classification since the annotations only include two labels: OOC and Not OOC. For the general fact-checking task, we evaluate models under a three-class setting, with labels True, False, and Unproven. For both evaluation settings, we report the following metrics: Weighted F1 score, accuracy, weighted recall, and weighted precision.
V-C2 Veracity Prediction Result
| Dataset | Task | Model | F1 | Acc | Rec | P |
| NewsCLIPpings | OOC | CCN* | – | 84.7 | – | – |
| Sniffer | 86.5 | 86.5 | 86.5 | 86.5 | ||
| LEMMA | 52.3 | 58.2 | 58.2 | 61.5 | ||
| DEFAME | 60.7 | 60.8 | 60.8 | 60.7 | ||
| AgentFact | 56.6 | 59.7 | 59.7 | 61.8 | ||
| Mocheg | Fact-Check | LEMMA | 28.0 | 37.5 | 37.5 | 40.6 |
| DEFAME | 44.9 | 46.0 | 46.0 | 47.3 | ||
| AgentFact | 51.9 | 48.2 | 48.2 | 49.6 | ||
| ClaimReview | Fact-Check | LEMMA | 16.5 | 14.5 | 14.5 | 87.3 |
| DEFAME | 59.2 | 52.1 | 52.1 | 81.7 | ||
| AgentFact | 69.2 | 67.0 | 67.0 | 74.6 | ||
| RW-Post | Fact-Check | LEMMA | 45.6 | 42.1 | 42.1 | 78.4 |
| DEFAME | 51.3 | 46.5 | 46.5 | 80.1 | ||
| AgentFact | 65.5 | 57.8 | 57.8 | 79.1 |
Results for CCN are reported as provided in the original paper, since access to the full dataset required for retraining is currently unavailable.
Table III presents the model performance on both the multimodal fact-checking task and the out-of-context (OOC) detection task.
3-class Fact-Checking Task
Across all three fact-checking datasets, AgentFact consistently achieves the best overall performance. On ClaimReview2024+ and RW-Post, which requires multimodal verification, and contains complex, real-world cases, AgentFact delivers the highest F1, accuracy, and recall, indicating its robustness in multi-step reasoning and evidence integration. Although its precision is slightly lower than DEFAME, the notable gains in recall across all three datasets suggest that AgentFact is more effective at identifying true misinformation cases.
On Mocheg (a purely textual input setting because the post image context is not provided), AgentFact substantially improves over LEMMA and DEFAME, demonstrating strong capability in handling challenging misinformation instances even without multimodal signals.
These results collectively show that AgentFact’s multi-agent design, reliability-aware evidence filtering, and iterative reasoning contribute to substantial improvements in fact-checking performance across diverse scenarios.
Out Of Context (OOC) Task
CCN and Sniffer serve as OOC-specific detectors, whereas LEMMA, DEFAME, and AgentFact are general-purpose fact-checking models. Sniffer achieves the highest accuracy (88.4), as expected for a specialized OOC system. AgentFact (56.6 F1 / 59.7 Acc) surpasses LEMMA and is competitive with DEFAME, demonstrating that multi-agent reasoning enables meaningful generalization even in settings outside fact-checking.
Overall Findings.
Across all datasets, AgentFact delivers the strongest performance among general-purpose fact-checking models. It improves fact-checking accuracy through multi-agent collaboration, evidence reliability assessment, and iterative reasoning. Moreover, its competitive performance on an OOC dataset demonstrates encouraging generalization ability, despite the task mismatch. These results validate the design principles behind AgentFact and highlight its robustness across diverse real-world misinformation scenarios.
V-D Performance on Explainability
We evaluate models’ explanation quality through human assessment, following recent LLM-based fact-checking studies [kim2024can, zhang2023towards]. We sample 40 instances from the RW-Post test set for annotation, using clear scoring criteria and detailed guidelines. Each annotator receives the outputs from AgentFact, LEMMA, and the ground-truth (GT) rationale, displayed in a unified format and randomized order with source identities fully blinded.444For each instance, the three outputs are randomly permuted and assigned anonymized identifiers (1–3), preventing annotators from inferring which source produced which explanation.
Annotation Scheme.
Annotators evaluate each output along three dimensions: (1) Reasoning Hallucination (unsupported or logically invalid inference), (2) Evidence Hallucination (misrepresented or fabricated evidence), and (3) Label Justification (whether the predicted label is well-supported). Each dimension is scored on a three-level ordinal scale (0–2), where lower scores indicate fewer errors and higher faithfulness. For example, Reasoning Hallucination is rated as: 0 = no hallucination (fully grounded), 1 = mild over-interpretation, and 2 = severe fabricated or unsupported reasoning. For presentation, we convert all scores to a normalized 0–1 range using
so that higher values consistently indicate better performance. Full annotation guidelines and the interface are provided in the Appendix -E.
Result Analysis.
We report the average of the three annotators’ scores in Table IV. As shown in Table IV, ground-truth rationales achieve near-perfect scores across all criteria (0.94–1.00), demonstrating their high faithfulness and reliability. This confirms that our LLM-assisted annotation pipeline provides a strong upper bound for explanation quality.
Differences between AgentFact and LEMMA are statistically significant across all three criteria. We apply the Wilcoxon signed-rank test to the paired differences for each instance and report the mean paired difference () in the table. Across the 40 aligned instances, the test yields for Reasoning (), Evidence (), and Label Justification (). We additionally compute 95% bootstrap confidence intervals over the paired differences (e.g., Reasoning: ), all of which exclude zero, confirming the robustness of the improvements.
| Criteria | GT | LEMMA | AgentFact | p-value | |
| Reasoning ↑ | 0.969 | 0.445 | 0.767 | ||
| Evidence ↑ | 1.000 | 0.796 | 0.946 | ||
| Label ↑ | 0.937 | 0.429 | 0.758 |
V-E Ablation Study of Agents
Table V reports the ablation results across the four agent groups. Overall, the full AgentFact system achieves the best performance, demonstrating the complementary roles of the planning, text retrieval, image retrieval and reasoning agents.
Effect of Removing the Strategy Planning Agent (w/o Agent-SP)
In this variant, we remove the planning module entirely. No reasoning plan or plan-derived search list is produced; instead, the query generation component receives only the claim and its context as input when constructing search queries. Removing Agent-SP results in moderate decreases in weighted F1 and accuracy. This indicates that structured verification plans help the system break down complex claims and guide focused evidence retrieval. Without such planning, the system becomes more reactive and less efficient in identifying the correct evidence, leading to reduced classification performance.
Effect of Removing the Temporary Reasoning Agent (w/o Agent-R-T)
Ablating the temporary reasoning agent leads to decreases in macro F1 and recall. This confirms that early-stage reasoning plays an important role in filtering noisy evidence and identifying which retrieved information is actually useful for the final decision. Without Agent-R-T, irrelevant or weakly related evidence is more likely to propagate into later stages, mildly degrading performance. Note that we do not report a variant removing all reasoning agents, as the reasoning module is structurally required to convert retrieved evidence into a final veracity decision; removing it would leave the system unable to produce meaningful predictions.
Effect of Removing the Image Retrieval and Analysis Agent (w/o Agent-IR)
Removing multimodal verification capabilities produces a clear drop in weighted F1 and recall. This highlights the importance of visual evidence—such as detecting tampering or miscaptioning—in verifying multimodal claims. Agent-IR provides complementary signals that textual evidence alone cannot capture, and its removal makes the system less robust to visually grounded misinformation.
Effect of Removing the Text Retrieval and Validation Agent (w/o Agent-TR)
The most substantial performance degradation occurs when text retrieval and validation are removed. Weighted F1 dramatically decreases, and accuracy drops sharply. These results indicate that high-quality textual evidence remains the primary foundation for fact-checking, and that Agent-TR’s query generation, filtering, and reliability-aware summarization are essential for accurate veracity prediction.
Overall Findings.
Each agent group contributes meaningfully to the success of AgentFact. Textual evidence retrieval (Agent-TR) is the most critical component, while multimodal signals (Agent-IR), structured planning (Agent-SP), and intermediate reasoning (Agent-R-T) provide additional improvements that collectively enhance robustness and reasoning quality. These findings validate the design of AgentFact’s multi-agent architecture and underscore the importance of evidence structuring and iterative reasoning in factual verification.
| Model | F1 (weighted) | F1 (macro) | Acc | P | Rec |
| w/o Agent SP | 0.593 | 0.470 | 0.527 | 0.766 | 0.527 |
| w/o Agent-IR | 0.603 | 0.468 | 0.534 | 0.775 | 0.534 |
| w/o Agent-TR | 0.469 | 0.328 | 0.373 | 0.763 | 0.373 |
| w/o Agent-R-T | 0.611 | 0.473 | 0.541 | 0.775 | 0.541 |
| Full System | 0.629 | 0.489 | 0.570 | 0.770 | 0.570 |
We provide additional analysis on model performance across varying levels of sample difficulty in the Appendix -D.
V-F Case study
From these two examples, we observe that AgentFact is capable of producing coherent reasoning and structured key points supported by multimodal evidence from generally reliable sources. However, a closer comparison with the ground truth reasoning reveals several notable shortcomings.
A common failure pattern highlighted by these cases is the model’s inability to retrieve accurate image contextual evidence, which is critical for effective multimodal fact-checking. While the retrieved evidence may appear relevant at a surface level, it often fails to capture the true historical or social context of the image.
Another limitation is that, although current agent-based methods can perform multi-step reasoning and aggregate information from multiple sources, they frequently lack a deep understanding of real-world claim context compared to human fact-checkers. In particular, miscaption detection is often reduced to verifying textual consistency between reverse image search results and the claim, without critically assessing the semantic relevance or informational quality of the retrieved evidence. For example, some retrieved texts consist primarily of commercial or non-contextual content (e.g., advertisements), which should not be treated as valid contextual evidence.
In contrast, human annotators demonstrate stronger contextual reasoning and evidence selection abilities, enabling more accurate evidence retrieval and, consequently, more reliable fact-checking judgments.
Claim:
A photograph genuinely shows singers Bono, Paul McCartney, and
Freddie Mercury performing together at the Live Aid benefit
concert in 1985.
| Aspect | Details |
| Label | Model: False GT: True |
| Reasoning |
Model:
Although Bono, Paul McCartney, and Freddie Mercury all appeared at
the Live Aid concert on July 13, 1985, the model concludes that
there is no explicit textual evidence confirming that the three
artists performed together on the same stage.
GT: The photograph is authentic and originates from the 1985 Live Aid concert. Multimodal evidence confirms that Bono, Paul McCartney, and Freddie Mercury appeared together during the same on-stage segment. |
| Key Points |
Model:
(1) All three artists individually participated in Live Aid.
(2) No official textual report explicitly documents a joint
performance.
GT: (1) The image is verified as originating from Live Aid (1985). (2) Video footage captures on-stage interaction between Bono and Freddie Mercury. (3) Archival photographs show Paul McCartney appearing alongside them. |
| Evidence |
Model Evidence:
[1] Text from the-paulmccartney-project.com: Paul McCartney performed at Live Aid on July 13, 1985. [2] Text from rollingstone.com: U2 (with Bono) performed at Live Aid on July 13, 1985. GT Evidence: [1] Video from YouTube: Bono and Freddie Mercury singing into the same microphone on stage. [2] Image/Text from Getty Images: Photographs depicting Bono, Paul McCartney, and Freddie Mercury appearing together during the concert. |
| Comparison | The model rejects the claim due to the absence of explicit textual documentation for a three-person performance, whereas the ground truth confirms the claim by integrating complementary multimodal evidence that correctly attributes the photograph to the Live Aid concert scene. |
Claim:
A photograph shows 26 corrupt politicians shortly before they were executed
in China.
| Aspect | Details |
| Label | Model: False GT: False |
| Reasoning |
Model:
The model concludes that the photograph is entirely miscaptioned, assigning
a 100% misrepresentation rate and asserting that the image is unrelated to
the alleged execution of corrupt politicians. No credible evidence is found
to support the claim.
GT: Fact-checking analysis confirms that the photograph was taken in April 2004 during a sentencing rally in Wenzhou, China, and depicts convicted criminals rather than politicians. |
| Key Points |
Model:
(1) The image is assessed as 100% miscaptioned.
(2) No reliable sources validate the execution claim.
GT: (1) The photograph originates from a 2004 sentencing rally in Wenzhou. (2) The individuals shown are convicted criminals, not politicians. (3) The claim misrepresents both context and identities. |
| Evidence |
Model Evidence:
[1] Image from Getty Images: search results for death row inmates (miscaption score: 100%). [2] Text from Human Rights Watch: background information on China’s anti-corruption campaign. GT Evidence: [1] Image from Snopes: photograph identified as a sentencing rally in Wenzhou (2004). [2] Text from Getty Images: original caption confirming the event as a sentencing rally involving convicted criminals. |
| Comparison | While the model correctly rejects the claim, it relies on a generic miscaption assessment and background political context. In contrast, the ground truth precisely establishes the photograph’s provenance and historical setting, offering a concrete explanation for why the claim is misleading. |
VI Conclusion
In this work, we introduce a high-quality dataset for explainable multimodal fact-checking, with rich annotations covering text-image claims, claim referenced posts, reasoning and aligned evidence. Additionally, we propose AgentFact, a novel agent framework that emulates human fact-checking by coordinating specialized agents for strategy-based planing, evidence analysis and reasoning, enabling interpretable and context-aware verification.
Despite its dedicated design, our approach still has several limitations that warrant further investigation in future work.
First, as discussed ine0989389@u.nus.edu the case study section, current agent-based fact-checking methods still fall short in understanding real-world contextual nuances of claims and evidence, which is crucial for more intelligent and reliable fact-checking.
Second, processing multiple pieces of evidence and images incurs considerable computational cost, which limits scalability. Exploring open-source models and lightweight deployment strategies is practically beneficial for reducing costs and supporting accessible fact-checking.
Third, the current memory design is not well suited for long-term reasoning across complex verification processes. Incorporating memory mechanisms tailored for long-term reasoning, such as experience-based memory and reflection modules, may further improve agent performance by enabling learning from past verification cases.
Finally, although we take care to avoid data leakage during evidence retrieval, some factual content may already be present in the pretraining data of large language models. Regular dataset updates and continual data curation are therefore necessary to ensure the freshness and reliability of fact-checking benchmarks.
DANNI XU is currently a Ph.D. student with the School of Computing, National University of Singapore. She received the B.S. degree in computer science from Wuhan University, Wuhan, China, in 2019, and the M.S. degree in computer science from the same university in 2022. Her research interests include misinformation detection, multimodal learning, and human behavior analysis. |
SHAOJING FAN is currently a Senior Lecturer in the Department of Electrical and Computer Engineering (ECE), National University of Singapore (NUS). Before joining ECE, she was a Senior Research Fellow in the School of Computing at NUS. Prior to that, she worked as a Senior Research Engineer at the Institute for Infocomm Research (I2R), part of Singapore’s Agency for Science, Technology, and Research (A*STAR). Dr. Fan received her B.E. and M.E. degrees from South China University of Technology in 2001 and 2004, respectively. She earned her D.Phil. from Ningbo University, with the majority of her research conducted at the I2R, A*STAR, between 2010 and 2015. Dr. Fan’s research interests include computer vision, cognitive vision, computational social science, and experimental psychology. Her work has been published in leading international journals and conferences, including TPAMI, CVPR, AAAI, and SIGGRAPH Asia. She is a member of the Association for Computing Machinery (ACM) and serves as a reviewer for several top-tier journals and conferences, such as TPAMI, CVPR, AAAI, ICLR, ICML, and NeurIPS. |
HARRY CHENG is a Research Fellow at the National University of Singapore, Singapore. He received his Ph.D. degree from Shandong University, China. He has authored or coauthored several papers in top conferences and journals, including NeurIPS, ICCV, ACM MM, and IEEE TMM, and he serves as a regular reviewer for several journals such as IEEE TPAMI, IEEE TIP, IEEE TIFS, IEEE TKDE, IEEE TMM, and ACM ToMM. |
MOHAN KANKANHALLI (Fellow, IEEE) is Provost’s Chair Professor of Computer Science at the National University of Singapore (NUS) and the Deputy Executive Chairman of AI Singapore. He is also the Director of NUS AI Institute, where he leads initiatives on multimodal models and trustworthy machine learning. Mohan obtained his BTech from IIT Kharagpur and MS & PhD from the Rensselaer Polytechnic Institute. Mohan’s research interests are in Multimodal Computing, Computer Vision and Trustworthy AI. Mohan was a member of World Economic Forum’s 2023-2024 Global Future Council on the Future of Artificial Intelligence. He is a member of ACM’s Global Technology Policy Council. Mohan is a Fellow of IEEE, IAPR and ACM. |
-A Dataset
Table S1 and Table S2 present the prompts used to filter samples in the RW-Post dataset, aiming to reduce label leakage from images and to exclude samples where only the text contributes to veracity assessment.
-B Prompts of Agents in AgentFact
-C Experimental Details: Fact-Checking Domain Filtering
To mitigate potential label leakage during evidence retrieval—particularly when external search engines return fact-checking articles that explicitly state the veracity of the claim—we applied a domain filtering strategy.
We excluded any evidence items whose source URLs contained substrings associated with known fact-checking organizations or services. This ensures that the model cannot trivially infer labels by relying on already-verified ground truth statements from professional sources. The filtering was implemented by checking whether the domain name or URL of the retrieved evidence contains any of the following substrings, which are associated with known fact-checking organizations or services. For brevity, a representative list is shown below; the complete list is available in our code repository.
snopes
politifact
factcheck
truthorfiction
hoax-slayer
eadstories
opensecrets
fullfact
checkyourfact
realitycheck
fact-check
This filter was applied consistently to all models that utilized external search (e.g., LEMMA, DEFAME, AgentFact) during evidence retrieval. By removing these high-authority verification sources, we ensure that the model’s performance more accurately reflects its ability to reason over raw evidence, rather than memorize or match against curated labels.
-D Experimental Results: Impact of Different Difficulty Levels.
To analyze model robustness under varying verification difficulty, we group claims based on the number of evidences referenced in their corresponding human-written fact-check articles. Importantly, these evidences are not provided to the model during inference. Instead, the evidence count is used solely as a proxy for the inherent difficulty of verifying a claim, reflecting the amount of effort required by human fact-checkers.
Figure S1 illustrates model performance across different difficulty levels for both binary (true vs. not_true) and three-class (true / false / unproven) verification settings. For binary evaluation, we collapse the original three-class labels into a two-class setting by treating true as the positive class and grouping all remaining labels (i.e., false and unproven) into a single negative class. This formulation reflects a practical fact-checking scenario where the primary objective is to determine whether a claim can be verified as true, while claims that are false or lack sufficient supporting evidence are both treated as not true.
Across both settings, we observe a clear non-monotonic performance trend with respect to claim difficulty. In particular, claims of moderate difficulty—those requiring 4 to 6 evidences in human fact-check articles—consistently yield the lowest Macro-F1 scores. This phenomenon is observed in both binary and three-class evaluations, indicating that such claims pose the greatest challenge to the model.
Interestingly, claims associated with very high evidence redundancy (9+ evidences) achieve the strongest performance in both evaluation settings. Despite being more complex in terms of information volume, these claims often correspond to highly visible events with strong cross-source consensus, which may facilitate model verification through information redundancy.
Overall, these results indicate that claim difficulty, as approximated by groundtruth evidence count, has a substantial impact on model robustness. Rather than exhibiting monotonic degradation, model performance follows a valley-shaped pattern, with moderate-difficulty claims being the most challenging. This highlights the importance of developing fact-checking models that are robust to fragmented and ambiguous evidence. We do not claim statistical significance for these differences, as the number of high-difficulty cases is relatively small.
-E Human evaluation
Figure S2 shows the interface used by human evaluators to score the outputs of different models, including the ground-truth results. The results of the three models (including the ground-truth) are randomly ordered for each sample, and the annotators are blind to which model produced each result.
The following guideline was given to human evaluators, outlining the definitions of all evaluation metrics and the corresponding rating criteria.
Task Definition
The goal of the evaluator (the annotator) is to assess the outputs produced by fact verification models. Each model aims to determine the authenticity of a given claim and provide corresponding explanations.
Each claim originates from a social media post. The post text and image are provided as contextual information to assist understanding, but the final evaluation focuses strictly on the claim itself.
Evaluation Interface Overview
In the annotation interface, each claim is presented together with its textual content, image context, and outputs generated by three different models. Each model output is shown as an independent card containing:
Authenticity judgment
Reasoning and explanation
Cited evidence
After finishing the evaluation for one claim, the results should be exported as a JSON file named according to the claim ID (e.g., 1.json, 2.json).
Each model output follows a structured JSON format:
"predicted_authenticity_class": "True",
"predicted_authenticity_reasons": "...",
"predicted_authenticity_key_points": [...],
"evidence_list": [...],
"image_analysis_result": [...],
"confidence_level": ""
Reasoning Hallucination
What to Evaluate
This dimension evaluates the factual soundness of the model’s reasoning. Specifically, check whether the reasoning provided in predicted_authenticity_reasons is directly supported by the evidence listed in evidence_list.
How to Judge
1.
Read the reasoning and key points.
2.
Identify factual claims such as numbers, dates, or causal statements.
3.
Verify whether these claims are supported by the cited evidence.
4.
If no concrete facts are invoked, evaluate logical coherence only.
Target Fields
predicted_authenticity_reasons
predicted_authenticity_key_points
| Score | Criteria |
| none | Reasoning is coherent and well-supported by evidence. |
| mild | Minor unsupported assumptions without affecting the main conclusion. |
| severe | Core reasoning is speculative or disconnected from evidence. |
Evidence Usage Hallucination
What to Evaluate
This dimension assesses whether the cited evidence is correctly used and accurately represented.
How to Judge
1.
Examine each item in evidence_list and image_analysis_result.
2.
Check whether the evidence supports the claims made.
3.
Identify any fabrication, misinterpretation, or irrelevance.
Target Fields
evidence_list
evidence_content_and_link
image_analysis_result
| Score | Criteria |
| none | Evidence is accurate, relevant, and supports the reasoning. |
| partial | Evidence is partially misused or weakly relevant. |
| full | Evidence is fabricated or contradicts the conclusion. |
Label Justification
What to Evaluate
This dimension examines whether the predicted authenticity label is justified.
How to Judge
1.
Identify the predicted label.
2.
Compare it with the reasoning and evidence.
3.
Check for overconfidence or contradiction.
Target Fields
predicted_authenticity_class
predicted_authenticity_reasons
| Score | Criteria |
| justified | Label is consistent with reasoning and evidence. |
| overconfident | Label is stronger than supported by evidence. |
| hallucinated | Label contradicts the reasoning or evidence. |
![[Uncaptioned image]](images/author/profile_danni.jpg)
![[Uncaptioned image]](images/author/profile_shaojing.jpg)
![[Uncaptioned image]](images/author/profile_xuanang.jpg)
![[Uncaptioned image]](images/author/profile_mohan_1.jpg)